any one help me in this prb


0.7g of Na^2CO^3.XH^2O is dissolved in 100mlH^20,20ml of which required 19.8 ml of 0.1N HCL.the value of x is

Grade:12

1 Answers

Askiitians Expert Bharath-IITD
14 years ago

Dear Jauneet,

We know that the hydrolysis of anhydrate sodium carbonate give 2 moles of sodium hydroxide.

So there is already x amount of water present in the sample so amount of water it contributes is x*18*density of water = 18x

We know that molarity= no of moles per litre of solution so the molecular weight of the given compound is (106+18x)gm/ mol and the total amount of water added is (100+18x) and we also know that basicity of the given compound is 2 two as it gives two moles of NaOH on hydrolysis. we also have the formula N1 * V1 =N2 * V2

=> 20*2*(0.7/(106+18x)) * (100/(100+18x)) = 0.1* 19.8

on solving this equatiopn we get x= 0.886 moles

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best  !!!

 


Regards,

Askiitians Experts

Adapa Bharath